OpenStudy (anonymous):

Two forces with magnitudes of 25 and 30 pounds act on an object at angles of 10° and 100° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

@Catlover5925 help me please?:)

OpenStudy (catlover5925):

I have no idea sorry, I will tag people to help @perl @hkhunny @iambatman @Ashleyisakitty @Kit_Kat_Nat_<3

OpenStudy (anonymous):

thanks anyway:)

OpenStudy (catlover5925):

@jordanloveangel @tkhunny

OpenStudy (jordanloveangel):

we can add the x and y components separately and use our knowledge of whether x and y are positive or negative in the quadrant in which it lies. x components: 25*cos(10) - 30*cos(80) = Fx y components: 25*sin(10) + 30*sin(80) = Fy Magnitude [absolute value]: F=∣∣Fx2+Fy2−−−−−−−−−√∣∣ and the direction of F is the angle define by: tan(θ)=FyFx

OpenStudy (jordanloveangel):

sin(x) = o/h cos(x) = a/h tan(x) = o/a o = opposite; a = adjacent; h = hypotenuse

OpenStudy (anonymous):

where does the 80 come from?

OpenStudy (jordanloveangel):

http://openstudy.com/study#/updates/515efe98e4b01c244f4a97c9

OpenStudy (jordanloveangel):

this might help u

OpenStudy (anonymous):

ok thank you:)

OpenStudy (jordanloveangel):

yah np wait did it help u